# Language models scale reliably with over-training and on downstream tasks

### 13 Mar 2024 v1

@@ 6 Limitations, future work, and conclusion
  • Hyperparameters. While our configurations are surprisingly amenable to reliable scaling across many training and testing distributions without further tuning, there is a need to further develop scaling laws that incorporate hyperparameters.
  • Scaling up. Validating the trends in this paper for even larger runs is a valuable direction.
  • Scaling down. Additionally, actualizing predictable scaling with even cheaper runs is important to make this area of research more accessible, especially for downstream error prediction.
  • Failure cases. While we present preliminary analysis of when scaling is unreliable, future work should develop an analytic theory explaining when scaling breaks down.
  • Post-training. It is common to employ supervised fine-tuning and reinforcement learning after pre-training, which we do not consider. Quantifying to what degree over-training the base model provides benefits after post-training is an open area of research.
@@ Conclusion.
We (i) show that the loss scaling behavior of models trained past compute-optimal, in the over-trained regime, is predictable and (ii) predict, via a proposed scaling law, the downstream average task performance of more expensive runs using smaller-scale proxies. We hope our work will inspire others to further examine the relationship between model training and downstream generalization. We also hope our testbed will make scaling research more accessible to researchers and practitioners alike.
@@ Appendix A Contributions
Project-specific infrastructure, plots, and analysis.
@@
Yair Carmon (original idea that “parallel lines” should show up in scaling plots), Samir Yitzhak Gadre (various derivations, empirical verification, related validation loss to average top-1 error as in Equation (5)), Reinhard Heckel (derived a scaling form based on Chinchilla Approach 3 [43], which appears in Equation (4)), Mitchell Wortsman (provided intuition about irreducible loss and why it is critical), Niklas Muennighoff (derived a scaling form based on Chinchilla Approach 3, similar to Equation (4)).
@@
Yair Carmon (co-lead), Achal Dave (co-lead), Alexandros G. Dimakis, Reinhard Heckel (co-lead), Gabriel Ilharco, Jenia Jitsev, Thomas Kollar, Niklas Muennighoff (co-lead), Ludwig Schmidt (co-lead), Shuran Song
@@ Appendix B Scaling-law derivations
This is equal to Equation (4), making the substitutions $\alpha_{C}=\alpha/2$, $a=A(1/6)^{-\alpha_{C}}$, $b=B(1/6)^{-\alpha_{C}}$, as noted in the main body.
@@
Appendix C Additional grid search details
@@
$n_{layers}$ $n_{heads}$ $d_{model}$ Number of parameters [B]
@@
$n_{layers}$ $n_{heads}$ $d_{model}$ Number of parameters [B]
@@
Appendix D Evaluation dataset details
@@
Downstream task LLM-foundry category Evaluation type Shots Samples Chance accuracy
@@
Appendix E Additional results
@@
As mentioned in Section 3.2 and Figure 4, we notice that models between 0.011B to 0.079B (i.e., $5.2\times 10^{16}$ to $5.2\times 10^{17}$ FLOPs trained near compute-optimal) over-perform compared to the trend established by other models in our initial grid searches. This results in a bump in the scaling plot. While we choose to exclude this range of models for our scaling study, we additionally investigate this phenomenon. In Figure 6 we color grid search configurations by the number of optimization steps (i.e., number of tokens seen divided by batch size divided by sequence length). For context, Figure 1 (left) in Kaplan et al. [49] also shows a bump; however, there the performance is worse than the general trend instead of better as in our work. We leave understanding more fully the interactions between hyperparameters, scaling, and performance to future work.
@@
Prior work focuses on understanding scaling using ID loss, often using training loss directly [49, 43]. Hence, we also consider Paloma [63] loss evaluation sets, which are designed to probe performance in specific domains. We use Paloma’s C4 [86, 25], RedPajama [110], and Falcon-RefinedWeb [80] splits to probe for ID validation loss. As seen in Figure 7, relative error is mostly low. Relative error is largest for the $N=1.4\text{B},M=640$ RedPajama run at 15.4%. Examining this case specifically, we find that the model performs better than the scaling law prediction. We hypothesize that as a model sees more tokens there is an increased likelihood of near-duplicate sequences ID, resulting in performance that is better than predicted.
@@
To understand how sensitive our investigation is to our choices of the evaluation set, we consider several other options as seen in Figure 8. We find that our prediction errors are fairly (i) low and (2) consistent for many choices of downstream evaluation sets.
Figure 8: Downstream evaluation set ablation for 6.9B parameter, 138B token runs. Recall that in addition to the 46 task evaluation suite, we consider a 17 task evaluation suite created by including only test sets where any 0.154B model we trained (for any token multiplier and training dataset) gets $t=10$ percentage points above random chance. Here we wish to understand how sensitive our results are to this choice $t$. (left) We see that relative prediction error is fairly low before a threshold of $t=35$. When too many tasks are excluded (i.e., $t>40$ relative error spikes. (right) A parallel view, showing how many tasks are removed as $t$ increases. 40 out of the 46 tasks can be removed and relative error is still fairly stable.
Scaling law fit Train set MMLU [38] OpenBook QA [66] HellaSwag [123] 17 eval 46 eval
Table 2 C4 [86, 25] 2.82% 16.80% 79.58% 0.14% 0.07%
Table 2 w/o 1.4B C4 [86, 25] 1.86% 96.16% 61.79% 0.42% 0.76%
Table 2 RedPajama [110] 0.12% 8.44% 25.73% 0.05% 2.10%
Table 2 w/o 1.4B RedPajama [110] 1.07% 7.56% 30.98% 10.64% 7.68%
Table 2 RefinedWeb [80] 0.77% 1.92% 81.96% 2.94% 2.76%
Table 2 w/o 1.4B RefinedWeb [80] 2.29% 6.79% 6.52% 15.79% 8.57%
Table 7: Downstream relative prediction error at 6.9B, 138B tokens, with and without the 1.4B data point. Recall in Table 2, we introduce a $N=1.4$B, $M=20$ run to get better downstream error predictions. Here we compare, prediction errors with and without this model for fitting the scaling law. Note that without the model (i.e., rows with “w/o 1.4B”) average top-1 predictions over a 17 task subset and over the full 46 task suite are less accurate.
@@
Figure 9: Scaling with small token multipliers. For smaller multipliers (e.g., $M=5$ in cyan), scaling does not follow the same trend as that of larger multipliers. Additionally, many token multiplies (e.g., $M\in\{10,20,40,80\}$) garner points close to the compute-optimal frontier.
@@
To further probe OOD performance, we measure the relative error of scaling laws fit to models trained on C4 and evaluated on Paloma’s 100 programming languages [63], Paloma’s Penn Tree Bank (PTB) split [64], and a German version of C4 [25]. Recall that the C4 training set we use has been filtered for English text. Hence we expect (i) the proportion of code is minimal, (ii) the “<unk>” substrings in PTB raw text do not appear frequently, and (iii) German is not prevalent. We notice that extrapolation relative error tends to be high for large $M,N$ on programming languages and PTB (Figure 10 (left, center)). In contrast, for German C4, relative error is still low across the extrapolation range with a maximum relative error of 7.6% at the $N=$1.4B, $M=80$ scale (Figure 10 (right)). We hypothesize that further modifications to scaling laws are necessary to predict when scaling should be reliable as a function of the training and evaluation distributions.
@@
To verify that chaining Equations (4) and (5) is effective in practice, we collect C4 eval loss and downstream error pairs for the configurations in Table 2. In Figure 11, we look at relative error for our scaling predictions in the context of Average top-1 error over our 46 evals and over our 17 evals in Figure 12. We again notice reliable scaling in interpolation and extrapolation regimes, suggesting the validity of our procedure to predict downstream average top-1 error.
@@
Figure 13 presents the correlation between downstream error vs. loss evaluated on different validation sets (C4, RedPajama, and RefinedWeb). Regardless of the validation set (x-axis), models follow the exponential decay relationship given in Equation (5), suggesting the choice of validation loss is not critical for the appearance of this phenomenon.
Figure 13: Correlation between average top-1 error and evaluation loss. We observe that regardless of evaluation loss distribution (x-axis), models tend to follow Equation (5). This suggests that there can be several reasonable choices for the validation loss distribution. Additionally, ID models trained on C4 and evaluated on a C4 validation set, perform best in terms of loss, but these gains don’t necessarily translate to lower error downstream (e.g., (left column)). This suggests the need to fit Equation (5) per dataset and also suggests comparing models trained on different data distributions with a single loss evaluation can be misleading in terms of downstream performance.
@@
Thus far we have looked at standard configurations from Table 2 to construct our scaling laws, mainly to demonstrate extrapolation to larger $N,M$. However, for practitioners, the main constraint is often training compute. Hence, we wish to understand the trade-offs between the amount of compute invested in creating a scaling law and the relative error of the resulting law in the over-trained regime. In Figure 14 (left), we see that as one increases the amount of compute, it is possible to get better fits with lower relative error. In Figure 14 (right), we see a similar trend as one increases the number of data points used to fit a scaling law. Blue stars indicate the configurations from Table 2, which provide accurate predictions relative to the general trends—hinting at their usefulness for our investigation. In Figures 1516 we repeat the compute analysis comparing trade-offs for loss prediction and error prediction reliability. We find that less compute is generally necessary to construct a loss scaling law that achieves the same relative error as that of an error prediction scaling law.
Figure 17: Scaling exponent vs. token multiplier. In Figure 2 we notice roughly parallel lines (i.e., roughly constant scaling exponent $\alpha_{C}$) in the $\log$-$\log$ plot of loss vs. compute, even as the token multiplier $M$ changes. Here we plot $\alpha_{C}$ vs. $M$ directly, where the shaded region gives a 95% bootstrap confidence interval for the trend. This view supports that $\alpha_{C}$ is relatively constant.
@@
Appendix F Additional related work

### 14 Jun 2024 v2

@@ 6 Limitations, future work, and conclusion
  • Hyperparameters. While our configurations are surprisingly amenable to reliable scaling across many training and testing distributions without further tuning, there is a need to develop scaling laws that do not require extensive hyperparameter sweeps.
  • Scaling up. Validating the trends in this paper for even larger runs is a valuable direction. Additionally, repeating our setup for models that achieve non-trivial performance on harder evaluations like MMLU is left to future work.
  • Scaling down. Actualizing predictable scaling with even cheaper runs is important to make this area of research more accessible, especially for downstream error prediction.
  • Failure cases. While we present a preliminary analysis of when scaling is unreliable, future work should investigate conditions under which scaling breaks down.
  • Post-training. It is common to employ fine-tuning interventions after pre-training, which we do not consider. Quantifying to what degree over-training the base model provides benefits after post-training is an open area of research.
@@ Conclusion.
We show that the loss of over-trained models, trained past compute-optimality, is predictable. Furthermore, we propose and validate a scaling law relating loss to average downstream task performance. We hope our work will inspire others to further examine the relationship between model training and downstream generalization. Our testbed will be made publicly available, and we hope it will make scaling research more accessible to researchers and practitioners alike.
@@ Appendix A Contributions
Project-specific planning, infrastructure, plots, and analysis.
@@
Yair Carmon (original idea that “parallel lines” should show up in scaling plots), Samir Yitzhak Gadre (various derivations, empirical verification, related validation loss to average top-1 error as in Equation (5)), Reinhard Heckel (derived a scaling form based on Chinchilla Approach 3 [45], which appears in Equation (4)), Niklas Muennighoff (derived a scaling form based on Chinchilla Approach 3, similar to Equation (4)), Mitchell Wortsman (provided intuition about irreducible loss and why it is critical).
@@
Yair Carmon (co-lead), Achal Dave (co-lead), Alexandros G. Dimakis, Reinhard Heckel (co-lead), Gabriel Ilharco, Jenia Jitsev, Pang Wei Koh, Thomas Kollar, Niklas Muennighoff (co-lead), Ludwig Schmidt (co-lead), Luca Soldaini, Shuran Song
@@ Appendix B Scaling-law derivations
This is equal to Equation (4), making the substitutions $\eta=\alpha/2$, $a=A(1/6)^{-\eta}$, $b=B(1/6)^{-\eta}$, as noted in the main body.
@@
Appendix C Additional training details
Architecture.
As stated in the main paper, we train transformers [116], based on auto-regressive, decoder-only, pre-normalization architectures like GPT-2 [85] and LLaMA [113]. We adopt OpenLM [39] for modeling, which utilizes PyTorch [80, 6], xformers [54], triton [75], FlashAttention [24], FSDP [130], and bfloat16 automatic mixed precision. Like LLaMA, we omit bias terms, but replace RMSNorm [128] with LayerNorm [8], which has readily available fused implementations. Following Wortsman et al. [123], we apply qk-LayerNorm [25], which adds robustness to otherwise poor hyperparameter choices (e.g., learning rate). We use SwiGLU [102] activations and depth-scaled initialization [129]. We use a sequence length of 2,048, rotary positional embeddings [106], and the GPT-NeoX-20B tokenizer [15], which yields a vocabulary size of 50k. We do not use weight tying [84, 46]. We sample without replacement during training and employ sequence packing without attention masking. We separate documents in our training corpora with end-of-text tokens.
Objectives and optimization.
[…]
Appendix D Additional grid search details
Final model configurations.
We present our final hyperparameters in Table 3.
Table 3: Main models and hyperparameters used in our investigation. Models have number of parameters $N$, with number of layers $n_{\text{layers}}$, number of attention heads $n_{\text{heads}}$, model width $d_{\text{model}}$, and width per attention head $d_{\text{head}}$. Batch sizes are global and in units of sequences. Each sequence has 2,048 tokens. A100 GPU hours are at $M=20$, which are near compute-optimal runs. For the 1.4B scale, a batch size of 256 performs slightly better than 512.
[…]
@@ Appendix D Additional grid search details
$n_{layers}$ $n_{heads}$ $d_{model}$ Number of
parameters [B]
@@
$n_{layers}$ $n_{heads}$ $d_{model}$ Number of
parameters [B]
@@
Appendix E Evaluation dataset details
@@
Downstream task LLM-foundry category Evaluation type Shots Samples Baseline
@@
Recall that we use a subset of 17 of these evaluations that give signal (are above random chance) for the compute range we consider. See Appendix F, where we ablate over the 17 subset design choice by including more and less evaluations.
Appendix F Additional results
Scaling law fits.
We present specific coefficients for our fits in Table 6.
Table 6: Scaling law fit parameters. Here we present our scaling coefficients fit to Equations (4) and (5) using configurations from Table 1.
Training dataset Fit for Equation (4): $L(C,M)=$ Fit for Equation (5): $\textsf{Err}(L)=$
$E+(a\cdot M^{\eta}+b\cdot M^{-\eta})C^{\eta}$ $\epsilon-k\cdot\exp{(-\gamma L)}$
C4 [88, 27] $1.51+\left(141\cdot M^{0.121}+190\cdot M^{-0.121}\right)C^{-0.121}$ $0.850-2.08\cdot\exp{(-0.756\cdot L)}$
RedPajama [112] $1.84+\left(212\cdot M^{0.136}+367\cdot M^{-0.136}\right)C^{-0.136}$ $0.857-2.21\cdot\exp{(-0.715\cdot L)}$
RefinedWeb [82] $1.73+\left(157\cdot M^{0.127}+246\cdot M^{-0.127}\right)C^{-0.127}$ $0.865-2.21\cdot\exp{(-0.707\cdot L)}$
Small-scale experiments can predict model rank order.
We expect to be able to rank hypothetical models based on their predicted performance, which is useful when deciding what large-scale runs to train. To verify, we rank 9 testbed models with $N\geq 1.4\text{B}$ by ground-truth top-1 error and by estimated top-1 error. We find high rank correlation of 0.88 for the 17-task split.
@@ Appendix F Additional results
As mentioned in Section 3.3 and Figure 4, we notice that models between 0.011B to 0.079B (i.e., $5.2\times 10^{16}$ to $5.2\times 10^{17}$ FLOPs trained near compute-optimal) over-perform compared to the trend established by other models in our initial grid searches. This results in a bump in the scaling plot. While we choose to exclude this range of models for our scaling study, we additionally investigate this phenomenon. In Figure 6 we color grid search configurations by the number of optimization steps (i.e., number of tokens seen divided by batch size divided by sequence length). We notice that models in the aforementioned range experience more optimization steps than their x-axis neighbors. For context, Figure 1 (left) in Kaplan et al. [51] also shows a bump; however, there the performance is worse than the general trend instead of better as in our work. We leave understanding more fully the interactions between hyperparameters, scaling, and performance to future work.
@@
Prior work focuses on understanding scaling using ID loss, often using training loss directly [51, 45]. Hence, we also consider Paloma [65] loss evaluation sets, which are designed to probe performance in specific domains. We use Paloma’s C4 [88, 27], RedPajama [112], and Falcon-RefinedWeb [82] splits to probe for ID loss. As seen in Figure 7, relative error is mostly low. Relative error is largest for the $N=1.4\text{B},M=640$ RedPajama run at 15.4%. Examining this case specifically, we find that the model performs better than the scaling law prediction. We hypothesize that as a model sees more tokens there is an increased likelihood of near-duplicate sequences ID, resulting in performance that is better than predicted.
@@
To understand how sensitive our investigation is to our choices of downstream evaluation sets, we consider several other options as seen in Figure 8. We find that our prediction errors are fairly (i) low and (ii) consistent for many choices of downstream evaluation sets including the whole suite of 46 evaluations.
Figure 8: Downstream evaluation set ablation for 6.9B parameter, 138B token runs. Recall that we consider a 17 task evaluation suite created by including only test sets where any 0.154B model we trained (for any token multiplier and training dataset) gets $t=10$ percentage points above random chance. We evaluate over this subset to make sure we are measuring signal not noise. Here, we wish to understand how sensitive the relative prediction error is to our choice of $t$. (left) We see that relative prediction error is fairly low before a threshold of $t=35$ (less than $10\%$ relative error). When too many tasks are excluded (i.e., $t\geq 40$) relative error spikes. Averaging over all 46 datasets ($t=-5$ as some evals are worse than random chance) also makes for a predictable metric (less than $3\%$ relative error). (right) A parallel view, showing how many tasks are removed as $t$ increases. 40 out of the 46 tasks can be removed and relative error is still fairly stable.
Table 7: Downstream relative prediction error at 6.9B, 138B tokens, with and without the 1.4B data point. Recall in Table 1, we introduce a $N=1.4$B, $M=20$ run to get better downstream error predictions. Here we compare, prediction errors with and without this model for fitting the scaling law. Note that without the model (i.e., rows with “w/o 1.4B”) average top-1 predictions, over the 17 tasks. are less accurate.
Scaling law fit Train set ARC-E LAMBADA OpenBook QA HellaSwag 17 eval
[23] [77] [68] [126]
Table 1 C4 [88, 27] 28.96% 15.01% 16.80% 79.58% 0.14%
Table 1 w/o 1.4B C4 [88, 27] 0.92% 2.04% 96.16% 61.79% 0.42%
Table 1 RedPajama [112] 5.21% 14.39% 8.44% 25.73% 0.05%
Table 1 w/o 1.4B RedPajama [112] 8.13% 11.07% 7.56% 30.98% 10.64%
Table 1 RefinedWeb [82] 26.06% 16.55% 1.92% 81.96% 2.94%
Table 1 w/o 1.4B RefinedWeb [82] 15.39% 6.26% 6.79% 6.52% 15.79%
@@
Figure 9: Scaling with small token multipliers. For smaller multipliers (e.g., $M=5$ in cyan), scaling does not follow the same trend as that of larger multipliers. Additionally, many token multipliers (e.g., $M\in\{10,20,40,80\}$) garner points close to the compute-optimal frontier.
@@
To further probe OOD performance, we measure the relative error of scaling laws fit to models trained on C4 and evaluated on Paloma’s 100 programming languages [65], Paloma’s Penn Tree Bank (PTB) split [66], and a German version of C4 [27]. Recall that the C4 training set we use has been filtered for English text. Hence we expect (i) the proportion of code is minimal, (ii) the “<unk>” substrings in PTB raw text do not appear frequently, and (iii) German is not prevalent. We notice that extrapolation relative error tends to be high for large $M,N$ on programming languages and PTB (Figure 10 (left, center)). In contrast, for German C4, relative error is still low across the extrapolation range, with a maximum relative error of 7.6% at the $N=$1.4B, $M=80$ scale (Figure 10 (right)). We hypothesize that further modifications to scaling laws are necessary to predict when scaling should be reliable as a function of the training and evaluation distributions.
@@
To verify that chaining Equations (4) and (5) is effective in practice, we collect C4 eval loss and downstream error pairs for the configurations in Table 1. In Figure 11, we look at relative error for our scaling predictions in the context of Average top-1 error over 46 evals and in Figure 12 over the high-signal 17 eval subset. We again notice reliable scaling in interpolation and extrapolation regimes, suggesting the validity of our procedure to predict downstream average top-1 error.
@@
Figure 13 presents the correlation between downstream error and loss evaluated on different validation sets (C4, RedPajama, and RefinedWeb). Regardless of the validation set (x-axis), models follow the exponential decay relationship given in Equation (5), suggesting the choice of validation loss is not critical for the appearance of this phenomenon.
Figure 13: Correlation between average top-1 error and evaluation loss. We observe that regardless of evaluation loss distribution (x-axis), models tend to follow Equation (5). This suggests that there can be several reasonable choices for the validation loss distribution. Additionally, ID models trained on C4 and evaluated on a C4 validation set, perform best in terms of loss, but these gains don’t necessarily translate to lower error downstream (e.g., (left column)). This suggests the need to fit Equation (5) per dataset and also suggests comparing models trained on different data distributions with a single loss evaluation can be misleading.
@@
Thus far we have looked at standard configurations from Table 1 to construct our scaling laws, mainly to demonstrate extrapolation to larger $N,M$. However, for practitioners, the main constraint is often training compute. Hence, we wish to understand the trade-offs between the amount of compute invested in creating a scaling law and the relative error of the resulting law in the over-trained regime. In Figure 14 (left), we see that as one increases the amount of compute, it is possible to get better fits with lower relative error. In Figure 14 (right), we see a similar trend as one increases the number of data points used to fit a scaling law. Blue stars indicate the configurations from Table 1, which provide accurate predictions relative to the general trends—hinting at their usefulness for our investigation. In Figures 15 and 16 we repeat the compute analysis comparing trade-offs for loss prediction and error prediction for our RedPajama 1.4B parameter, 900B token and 6.9B parameter, 138B token runs respectively. We find that less compute is generally necessary to construct a loss scaling law that achieves the same relative error as that of an error prediction scaling law.
Figure 17: Scaling exponent vs. token multiplier. In Figure 2, we notice roughly parallel lines (i.e., roughly constant scaling exponent $\eta$) in the $\log$-$\log$ plot of loss vs. compute, even as the token multiplier $M$ changes. Here we plot $\eta$ vs. $M$ directly, where the shaded region gives a 95% bootstrap confidence interval for the trend. This view supports that $\eta$ is relatively constant.
@@
On compute-optimal token multipliers.
We consider 20 tokens per parameter as close to compute-optimal for our experiments. Here we investigate, using different approaches, what the compute-optimal token multipliers are for each dataset—assuming one should scale number of parameter and training tokens equally as Hoffmann et al. [45] suggest.
Turning to Figure 9, we notice that there are many multipliers, between 10 and 80 that yield models close to the frontier. Hence, empirically, it appears choices within this range should be suitable for the optimal token multiplier.
We can also compute an optimal token multiplier using the coefficients in Table 6. Based on Hoffmann et al. [45]’s Equation (4) and the assumption that $\alpha=\beta$, we write,
$N^{*}(C)=G\left(\frac{C}{6}\right)^{\frac{1}{2}},D^{*}(C)=G^{-1}\left(\frac{C}{6}\right)^{\frac{1}{2}},G=\left(\frac{a}{b}\right)^{\frac{1}{4\eta}}.$ (9)
To compute $M^{*}=D^{*}/N^{*}$, we then have,
$M^{*}=\left(\frac{b}{a}\right)^{\frac{1}{2\eta}}.$ (10)
Using the values from Table 6 and Equation (10), we find $M^{*}_{\text{C4}}=3.36$, $M^{*}_{\text{RedPajama}}=7.42$, $M^{*}_{\text{RefinedWeb}}=5.85$, where the subscript gives the dataset name. These values conflict with the observation in Figure 9, which suggests $M=5$ is already too small to give points on the Pareto frontier. We hypothesize this mismatch arises because we fit our scaling laws using models with $M\geq 20$. Additionally, we hyperparamter-tune at $M=20$. As previously discussed, it is likely possible to find better hyperparameter configurations at $M=5$ with further hyperparameter tuning at this token multiplier.
Appendix G Additional related work